### v1

# An Adjoint Formulation of Energetic Particle Confinement

Abstract
An adjoint formulation of energetic particle confinement in axisymmetric geometry is derived and evaluated using a Physics-Informed Neural Network (PINN). The PINN estimates the escape time of energetic ions by solving an inhomogeneous adjoint of the drift kinetic equation with a Lorentz collision operator, yielding predictions of the escape time of fast ions in tokamak geometry due to direct ion orbit loss and collisional transport. This is the first time a PINN has been used to solve the drift kinetic equation in tokamak geometry, a challenging problem due to the large time scale separation present between the rapid transit time of energetic ions, and their slow collision time scale. It is shown that a careful and intentional design of a PINN is able to learn the escape time for the majority of the geometry considered, suggesting a path toward constructing a rapid surrogate for use in a broader optimization framework.
I Introduction
Energetic particles are a ubiquitous aspect of magnetic fusion plasmas. Such particles emerge as a result of ion heating schemes including neutral beam injection Hemsworth et al. (2008) and ion cyclotron resonance heating (ICRH) Adam (1987); Kazakov et al. (2017). For a burning deuterium-tritium plasma, the confinement of energetic alpha particles is critical for the sustainment of the fusion plasma. In addition, when such particles escape they can cause substantial damage to plasma-facing components Fasoli et al. (2007); Bonofiglo et al. (2024). Obtaining an understanding of the confinement time of such particles is thus vital to the operation of a magnetic fusion device.
In the present work, an initial step toward a framework through which the confinement time of an energetic ion or electron can be efficiently estimated across a broad range of parameter regimes is developed. While evaluating the confinement time of an energetic particle is often of greatest importance for non-axisymmetric magnetic fields arising from error fields Goldston et al. (1981), magnetic islands Carolipio et al. (2002) or energetic particle modes Heidbrink and White (2020) in a tokamak, or for a 3D stellarator equilibrium Galeev et al. (1969), this work will consider the simpler problem of an axisymmetric tokamak equilibrium. Our motivation for considering this simpler problem is largely to investigate the potential for evaluating metrics of energetic particle transport via the solution to an inhomogeneous adjoint problem. As described below, the inhomogeneous adjoint of the drift kinetic equation provides a direct means of computing the average confinement time of an ion or electron as a function of the particle’s initial phase space position Särkimäki (2020), a quantity we will refer to as the escape time. While this adjoint problem can be solved by a variety of numerical methods, we will consider a Physics-Informed Neural Network (PINN) Raissi et al. (2019) in the present work. Physics-Informed Neural Networks correspond to a machine learning framework capable of integrating both physics and data into the training of a neural network. A distinguishing feature of a PINN is that it may be extended to learn the parametric dependence of the solution of a PDE Sun et al. (2020). Noting that optimizing the transport properties of energetic particle orbits often requires exploring a broad range of magnetic field configurations, obtaining a rapid surrogate capable of computing quantities of interest such as the escape time of energetic particles provides a powerful tool for assessing energetic particle transport.
Our aim in the present paper is not to provide such a parametric solution, but to instead investigate the ability of a PINN to solve the drift kinetic equation for an energetic ion population. Specifically, while PINNs have previously been used to solve the relativistic Arnaud et al. (2024); McDevitt et al. (2025); Arnaud et al. (2025) and non-relativistic McDevitt and Tang (2024) Fokker-Planck equations, the large time scale separation between the rapid bounce or transit times and slow collision time of energetic particles in tokamak geometry, represents a far more challenging problem. Utilizing advanced optimization algorithms Urbán et al. (2025); Kiyani et al. (2025); Wang et al. (2025), we aim to show that PINNs are able to robustly resolve several features of this challenging system, even in the absence of data. As described further below, as the energy of the ions is increased, leading to a greater time scale separation between the fast bounce/transit motion and slow collision time of energetic ions, our implementation of a PINN struggles to quantitatively describe the escape time of the best confined ion orbits, though it does resolve the solution in the outer region of the tokamak. In future work, we will seek to overcome this limitation by directly utilizing small quantities of Monte Carlo data generated by a traditional particle based drift kinetic solver.
As an alternative means of evaluating the fast ion escape time, we introduce a new drift-kinetic solver based on the JAX framework. While in the present work this solver is only used to verify solutions from the PINN, future work will leverage particle data generated by this drift-kinetic solver to help accurately resolve the escape time of well confined ion orbits, along with facilitating the training of the PINN over a broad range of plasma parameters. This extended PINN framework, trained using both physics and particle data, would thus provide an efficient surrogate for fast ion confinement appropriate for use inside optimization loops.
The remainder of this paper is organized as follows. Section II derives an adjoint problem that will be solved in this paper. A brief overview of the physics-informed deep learning framework and JAX particle-based solver are described in Sec. III. The escape time of an energetic ion species is computed in Sec. IV. Conclusions and a discussion of future directions are given in Sec. V
II Adjoint of Steady State Drift Kinetic Equation
In this section we will define an adjoint problem of the inhomogeneous steady state drift kinetic equation. Our starting point will be to define the Green’s function $F_{s}\left(\mathbf{z};\mathbf{z}_{0}\right)$ for the steady state adjoint equation Helander and Sigmar (2002):
$\nabla\cdot\left(\mathbf{\dot{X}}F_{s}\right)+\frac{\partial}{\partial v_{\|}}\left(\dot{V}_{\|}F_{s}\right)-C_{s}\left(F_{s}\right)=\delta\left(\mathbf{z}-\mathbf{z}_{0}\right),$ (1a)
where $\mathbf{z}=\left(\mathbf{x},v_{\|},\mu\right)$ is the phase space position of the particle and the characteristic equations are given by:
$\mathbf{\dot{X}}=v_{\|}\bm{\hat{b}}+\mathbf{v}_{EB}+\mathbf{v}_{Ds},$ (1b)
$\dot{V}_{\|}=\frac{q_{s}}{m_{s}}\bm{\hat{b}}\cdot\mathbf{E}-\mu\bm{\hat{b}}\cdot\nabla B.$ (1c)
Here, the subscript $s$ is the species index, $\mu\equiv v^{2}_{\perp}/\left(2B\right)$ is the magnetic moment, $v_{\|}$ is the velocity parallel to the magnetic field, $C_{s}\left(F_{s}\right)$ is a collision operator, taken to be a Lorentz collision operator
$C_{s}\left(F_{s}\right)=\frac{\nu_{D}}{2}\frac{\partial}{\partial\xi}\left[\left(1-\xi^{2}\right)\frac{\partial F}{\partial\xi}\right]{\color[rgb]{1,0,0},}$ (2)
and we have defined the $\mathbf{E}\times\mathbf{B}$ and magnetic drifts as
$\mathbf{v}_{EB}=\frac{\mathbf{E}\times\mathbf{B}}{B^{2}},$ (3a)
$\mathbf{v}_{Ds}=\frac{\left(v^{2}_{\|}+\mu B\right)}{\omega_{cs}}\bm{\hat{b}}\times\nabla\ln B.$ (3b)
The magnetic drifts have been simplified by the approximation $\nabla\times\mathbf{\hat{b}}\approx\mathbf{\hat{b}}\times\nabla\ln B$, appropriate in the limit of a “vacuum plasma.” As boundary conditions we will enforce $F_{s}=0$ at the spatial boundaries whenever $U_{r}\equiv\mathbf{r}\cdot\mathbf{\dot{X}}<0$, where $\mathbf{\hat{r}}$ is the radial unit vector, and we have assumed a circular spatial boundary, but leave $F_{s}$ unconstrained for $U_{r}\equiv\mathbf{r}\cdot\mathbf{\dot{X}}>0$, i.e.
$F_{s}\left(r=r_{wall}\right)=\begin{cases}0,&U_{r}<0\\ \text{unconstrained},&U_{r}\geq 0\end{cases},$ (4)
where $r_{wall}$ is the location of the surrounding wall, taken as the plasma minor radius $a$. With these boundary conditions no flow of particles is allowed to enter the spatial domain, but particles are free to exit. The resulting Green’s function can thus be physically interpreted as the steady state particle distribution resulting from a unit source located at $\mathbf{z}=\mathbf{z}_{0}$.
To derive an adjoint problem for the escape time of an energetic particle, we begin by multiplying the left hand side of Eq. (1a) by the solution to the adjoint equation $T_{s}$, which after successive integrations by parts yields:
$\int d^{3}vd^{3}xT_{s}$ $\left[\nabla\cdot\left(\mathbf{\dot{X}}F_{s}\right)+\frac{\partial}{\partial v_{\|}}\left(\dot{V}_{\|}F_{s}\right)-C_{s}\left(F_{s}\right)\right]$
$=\int d^{3}vd^{3}xF_{s}\left[-\mathbf{\dot{X}}\cdot\nabla T_{s}-\mathbf{\dot{X}}\frac{\partial T_{s}}{\partial v_{\|}}-C^{*}_{s}\left(T_{s}\right)\right]$
$+\int d^{3}v\int d\Gamma\left.\left[U_{r}T_{s}F_{s}\right]\right|_{r=a}.$ (5)
Here, the first term on the right hand side (RHS) is the Green’s function times the adjoint of the steady state drift kinetic equation, and the second term on the RHS describes fluxes through the spatial boundary $\Gamma$. The boundary condition on the Green’s function $F_{s}$ [Eq. (4)] implies the spatial boundary term will vanish when $U_{r}<0$, where setting the boundary conditions on $T_{s}$ when $U_{r}\geq 0$, allows for distinct adjoint problems to be defined.
An adjoint problem for the escape time of an energetic particle can be derived by considering the inhomogeneous adjoint equation Liu et al. (2017)
$\mathbf{\dot{X}}\cdot\nabla T_{s}+\dot{V}_{\|}\frac{\partial T_{s}}{\partial v_{\|}}+C^{*}_{s}\left(T_{s}\right)=-1.$ (6)
Substituting Eq. (6) along with the steady state Green’s function defined by Eq. (1a) into the adjoint relation defined by Eq. (5), yields
$T_{s}\left(\mathbf{x}_{0},\mathbf{v}_{0}\right)$ $=N_{s}+\int d\Gamma\int d^{3}v\left.\left[U_{r}T_{s}F_{s}\right]\right|_{r_{wall}},$ (7)
where $\Gamma$ is the surface of the surrounding wall and $N_{s}\equiv\int d^{3}v\int d^{3}xF_{s}$ is the number of particles remaining in the domain due to the unit source on the right hand side of Eq. (6). At the spatial boundary we will take
$T_{s}\left(r=r_{wall}\right)=\begin{cases}\text{unconstrained},&U_{r}\leq 0\\ 0,&U_{r}>0\end{cases}.$ (8)
This boundary condition follows since a particle born on the spatial boundary with an outward directed velocity would be lost immediately, such that $T_{s}=0$, however, for an inward directed velocity this particle would not be immediately lost. For the boundary condition defined by Eq. (8), and noting that $F_{s}=0$ for $U_{r}<0$ [Eq. (4)], yields
$T\left(\mathbf{x}_{0},\mathbf{v}_{0}\right)$ $=N_{s}.$ (9)
Further noting that $N_{s}$ will scale directly with the confinement time of the particles, $T_{s}\left(\mathbf{x}_{0},\mathbf{v}_{0}\right)$ will act as a measure of how well particles injected at a given phase space location are confined on average, a quantity we will refer to as the escape time.
III Physics-Constrained Deep Learning Framework
The primary focus of the remainder of this work will be the solution of the adjoint problem derived in Sec. II. A brief overview of our implementation of a PINN is given in Sec. III.1, where the reader is referred to Refs. Cuomo et al. (2022); Wang et al. (2023); Toscano et al. (2025) for additional information on this rapidly expanding area. In addition, a newly developed GPU accelerated particle-based solver to the drift kinetic equation is described in Sec. III.2. For the present work, this particle-based solver will be used to validate the PINN’s solutions to the drift kinetic equation.
III.1 Physics-Informed Neural Networks
Physics-constrained machine learning methods Karniadakis et al. (2021) seek to incorporate physical constraints into the training of a neural network. Physics-informed neural networks correspond to a prominent example of this area. An attractive feature of PINNs is that they can be used to solve PDEs and ODEs directly in the absence of data van Milligen et al. (1995); Raissi et al. (2019), or they provide a mechanism through which physical constraints can be used to enrich sparse data sets Cai et al. (2021); Mathews et al. (2022); McDevitt et al. (2024). In its simplest form, the loss function of a PINN can be written as:
$\text{Loss}=\frac{1}{N_{data}}\sum^{N_{data}}_{i=1}\left[T_{i}-T\left(\mathbf{z}_{i};\bm{\lambda}_{i}\right)\right]^{2}+\frac{w_{pde}}{N_{PDE}}\sum^{N_{PDE}}_{i=1}\mathcal{R}^{2}\left(\mathbf{z}_{i};\bm{\lambda}_{i}\right),$ (10)
where $T$ is the quantity we seek to predict, the escape time for the current work, $\mathcal{R}$ is the residual of the governing PDE, $w_{PDE}$ is a scalar weight applied to the PDE bias of the loss, $\mathbf{z}=\left(\mathbf{x},\mathbf{v}\right)$ are the independent variables, and $\bm{\lambda}$ describes physical parameters, such as collisionality or inverse aspect ratio. The first term represents a bias due to data, where the data may be boundary conditions, or synthetic or experimental data. The second term is a bias due to the governing equations, which for the present work will be the inhomogeneous adjoint of the steady state drift kinetic equation. In the absence of synthetic or experimental data, the first term will only contain the boundary conditions. Successfully minimizing the loss will thus correspond to identifying a solution to the PDE while satisfying the boundary conditions. Unlike traditional PDE solvers, a parametric solution to a PDE can be straightforwardly obtained by training across a broad range of physics parameters $\bm{\lambda}$. Thus, while the training time of such a model may be substantially longer than a traditional PDE solver, once trained, the fast inference time of a PINN allows for fast online prediction time across a broad range of plasma conditions. This property can be used for the development of fast surrogate models, where some recent examples include runaway electron generation McDevitt (2023); Arnaud et al. (2025), MHD equilibrium Jang et al. (2024), plasma thrusters Luo et al. (2025), or non-thermal ion distributions McDevitt and Tang (2024).
A key element in obtaining accurate results from PINNs corresponds to (i) the identification of an appropriate neural network architecture, (ii) the phase space weighting of the residual, and (iii) the selection of an efficient optimization algorithm for minimizing the loss. Regarding item (i) we will employ a fully connected feedforward neural network, but with an ‘output layer’ that enforces positivity as a hard constraint. In so doing, this both prevents certain unphysical solutions from being predicted by the PINN, and perhaps more importantly, narrows the space of solutions that the optimizer searches for. Defining $T_{NN}$ to be the output of the hidden layers of the neural network, a simple transformation that enforces positivity is given by
$T_{s}=\exp\left(T_{NN}\right).$ (11)
In addition to enforcing positivity, the exponential dependence of $T_{s}$ on the output of the hidden layers $T_{NN}$ provides a natural means of capturing the large disparity in escape times for energetic ions. Specifically, ions born near the plasma core are expected to be well confined, where they will only be lost by the relatively slow process of collisional transport, in contrast to those born near the edge, which may be lost on a very short timescale.
An additional challenge that arises is ensuring that the PINN is able to robustly connect the boundary condition Eq. (8) applied at the edge of the tokamak, to the solution throughout the core plasma. Noting that we expect the escape time $T_{s}$ to be small near the edge, training points in this edge region can be given greater emphasis by weighting the residual of the PDE by:
$\mathcal{F}\left(\mathbf{z};\bm{\lambda}\right)=\frac{A}{A+T_{s}\left(\mathbf{z};\bm{\lambda}\right)},$ (12)
where $A$ is a constant, such that the loss can be written as:
$\text{Loss}=\frac{1}{N_{bdy}}\sum^{N_{bdy}}_{i=1}\left[T_{s,i}-T_{s}\left(\mathbf{z}_{i};\bm{\lambda}_{i}\right)\right]^{2}+\frac{w_{pde}}{N_{PDE}}\sum^{N_{PDE}}_{i=1}\left[\mathcal{F}\left(\mathbf{z}_{i};\bm{\lambda}_{i}\right)\mathcal{R}_{IH}\left(\mathbf{z}_{i};\bm{\lambda}_{i}\right)\right]^{2}.$ (13)
Here, the boundary conditions defined by Eq. (8) are enforced by the first term, $\mathcal{R}_{IH}$ is the residual of the inhomogeneous adjoint defined by Eq. (6), and the choice of $A$ sets how much emphasis is given to the edge region. In particular, noting that near the edge $T_{s}$ will be of the order of a transit time, since ions on loss orbits will directly escape from the plasma, if $A$ is chosen to satisfy $A\ll 1$, Eq. (12) will scale as $\propto 1/T_{s}$, such that the edge region will be heavily weighted, but the inner region of the plasma with $T_{s}\gg 1$, will receive minimal weighting. We thus expect this choice to allow for the outer region of the plasma to be accurately described, at the expense of poor accuracy in the inner region. In contrast, by choosing a larger value of $A$, improved accuracy for the inner region can be achieved, at the expense of having a looser connection to the edge boundary condition, which is vital for achieving a physically meaningful solution. We have found that values of $A$ between 10 and 100 allow for robust convergence of the PINN.
In addition, the relative weight of the PDE residual compared to the boundary condition is set by the scalar $w_{PDE}$ weighting the first term in Eq. (13). We have found $w_{PDE}=100$ provides relatively rapid convergence of the loss. Finally, since we anticipate the edge region to contain the most rapid variation of the escape time, we will pack a disproportionate number of training points into this region.
The last key component of our PINN implementation will be the choice of an effective optimization algorithm. While the ADAM optimizer Kingma and Ba (2014) is used ubiquitously across a range of machine learning applications, this first order optimization algorithm often is not able to achieve low losses for challenging PDEs. A common strategy is thus to use ADAM for the first phase of training, after which limited memory BFGS (L-BFGS) Liu and Nocedal (1989), a second order optimization strategy, is used to more tightly converge the loss. Motivated by Refs. Wang et al. (2025) and Urbán et al. (2025); Kiyani et al. (2025), we have adopted a different optimization strategy that we have found to yield substantially improved accuracy. The first optimizer that we apply corresponds to the quasi second order optimizer SOAP Vyas et al. (2024). This scalable optimizer exhibits similar computational performance as ADAM, but can yield losses of up to an order of magnitude smaller than those achievable by ADAM across a broad range of challenging applications of PINNs Wang et al. (2025). For the second stage of optimization, we utilize the quasi-Newton Self-Scaled Broyden (SSBroyden) method Al-Baali et al. (2014). Like L-BFGS, SSBroyden approximates the Hessian, but introduces additional degrees of freedom into the optimization algorithm that have yielded substantial improvements in training across a range of complex PDEs Urbán et al. (2025); Kiyani et al. (2025). While each training epoch is substantially more computationally demanding compared to ADAM or SOAP, we have found it results in substantially faster convergence of the loss for the adjoint problem treated in this work. The primary limitations of SSBroyden as an optimization routine is that the SciPy implementation used in the present work does not efficiently utilize GPUs or support mini-batching. These limitations substantially increase the training time of the SSBroyden phase of training, and limit the number of training points employed.
III.2 Jonta: A Particle Based Drift Kinetic Solver
Particle based drift kinetic simulations of fast ion orbits is computationally intensive White and Chance (1984); Hirvijoki et al. (2014), which requires moderate to substantial resources and efficient software. While this has predominately been done with Fortran or C++ libraries, with GPU acceleration facilitated by libraries such as Kokkos Trott et al. (2021), maintaining the support of these solvers can be cumbersome, due to its dependency on external backend libraries to effectively utilize high performance computing (HPC) hardware. Perhaps most importantly, given that PINNs are trained on GPUs and libraries with a large user base, this motives the development of a traditional drift kinetic solver that runs on GPUs and utilizes the same libraries that are used by PINNs, facilitating the operation of the broader framework that contains an interplay between the PINN and a drift kinetic solver.
With the aforementioned discussion, we present Just anOther fuNcTionAl pusher (JONTA), which is built on the PyTorch Paszke et al. (2019) and JAX Bradbury et al. (2018) libraries that run on GPU accelerated hardware. For the present work, we will utilize the JAX backend, and employ a four stage Runge Kutta (RK4) integration scheme, which is part of a broader suite containing differentiable integrator methods in the Optax library opt (2020). Guiding center equations are evolved with RK4, and collisions are implemented with a Monte-Carlo operator described below. Upon acceptance of the manuscript, the code will be uploaded to the public repository, where additional implementations can also be found for evolving charged particles in other fusion scenarios, such as runaway electrons in tokamak disruptions Breizman et al. (2019).
JONTA will be used to solve the guiding center equations defined by Eq. (1) together with a Monte Carlo collision operator describing pitch-angle scattering. The guiding center equations can be written as
$\dot{X}=\hat{\mathbf{x}}\cdot\dot{\mathbf{X}},$ (14a)
$\dot{Z}=\hat{\mathbf{z}}\cdot\dot{\mathbf{X}},$ (14b)
$\dot{v}=\frac{\partial v}{\partial v_{\|}}\frac{dv_{\|}}{dt}+\frac{\partial v}{\partial\mu}\frac{d\mu}{dt}+\frac{\partial v}{\partial\mathbf{X}}\cdot\frac{d\mathbf{X}}{dt},$ (14c)
$\dot{\xi}=\frac{\partial\xi}{\partial v_{\|}}\frac{dv_{\|}}{dt}+\frac{\partial\xi}{\partial\mu}\frac{d\mu}{dt}+\frac{\partial\xi}{\partial\mathbf{X}}\cdot\frac{d\mathbf{X}}{dt},$ (14d)
where we have chosen cylindrical coordinates $\left(R,Z\right)$, $X\equiv R-R_{0}$, and make the transformation from $\left(\mathbf{x},v_{\|},\mu\right)$ to $\left(\mathbf{x},\xi,v\right)$. Assuming an axisymmetric tokamak geometry with circular geometry and noting the relations $\xi^{2}=v^{2}_{\|}/\left(v^{2}_{\|}+2\mu B\right)$, and $v^{2}=v^{2}_{\|}+2\mu B$, Eq. (14) can be written as
$\dot{X}=\hat{\mathbf{x}}\cdot v_{\|}\bm{\hat{b}}=-\xi v\frac{Z}{qR_{0}},$ (15a)
$\dot{Z}=\hat{\mathbf{z}}\cdot v_{\|}\bm{\hat{b}}+\hat{\mathbf{z}}\cdot\mathbf{v}_{Ds}=\xi v\frac{X}{qR_{0}}-\frac{1}{2}\rho^{*}_{i}\left(1+\xi^{2}\right)v^{2}\frac{a}{R_{0}},$ (15b)
$\dot{v}=0,$ (15c)
$\dot{\xi}=\frac{\partial\xi}{\partial v_{\|}}\frac{dv_{\|}}{dt}+\frac{\partial\xi}{\partial\mathbf{X}}\cdot\frac{d\mathbf{X}}{dt}=-\frac{1}{2}\left(1-\xi^{2}\right)v\frac{Z}{qR_{0}}\frac{a}{R_{0}}\frac{B}{B_{0}}.$ (15d)
Here, we normalized time to $a/v_{Ti}$, where $a$ is the plasma minor radius taken to be equal to $r_{wall}$, speed has been normalized to $v_{Ti}$, space to $a$, and we have defined $\rho^{*}_{i}=\rho_{i}/a$, $\rho_{i}=v_{Ti}/\omega^{(0)}_{c}$, $\omega^{(0)}_{c}=ZeB_{0}/m_{i}$, $B_{0}$ is the magnetic field on axis, and $q$ is the safety factor. For convenience we have neglected the electric field, and assumed the limit of a small inverse aspect ratio and low-$\beta$ plasma. Specifically, the magnetic field was taken to have the form:
$\mathbf{B}=I\left(\psi\right)\nabla\varphi+\nabla\varphi\times\nabla\psi,$ (16)
$B_{\varphi}\left(r,\theta\right)=\frac{B_{0}}{1+\epsilon\cos\theta},$ (17)
$B_{\theta}\left(r,\theta\right)=\frac{B_{\theta}\left(r\right)}{1+\epsilon\cos\theta},$ (18)
with
$R=R_{0}+r\cos\theta,$ (19)
$q=\frac{rB_{0}}{R_{0}B_{\theta}\left(r\right)},$ (20)
and the poloidal flux function is related to the poloidal magnetic field by $d\psi/dr=R_{0}B_{\theta}\left(r\right)$. The magnitude of the total magnetic field can be conveniently written as
$\frac{B\left(r,\theta\right)}{B_{0}}=\frac{R_{0}}{R}\sqrt{1+\left(\frac{r}{qR_{0}}\right)^{2}}\approx\frac{R_{0}}{R},$
where in the last line we take $\left(a/R_{0}\right)^{2}\ll 1$.
Equations (15a) and (15b) describe the spatial evolution of the energetic ions due to parallel streaming and a vertical drift arising from the grad-B and curvature drifts. Our motivation for neglecting the electric field is that in this initial study we will assume stationary and homogeneous density, temperature and current profiles, such that no electric field is expected to arise. In the absence of the electric field, the collisionless velocity space dynamics are set by the mirror force, which only impacts the pitch of the ions such that the speed $v$ will be constant. While a more complete set of guiding center equations can be easily adopted, for this initial study, Eq. (15) will prove sufficient to capture several non-trivial aspects of the collisionless orbit of energetic ions.
Figure 1: Change of toroidal canonical momentum for 20 keV ions (panel a.) and 50 keV ions (panel b). Ten million deuterium ions were initialized randomly across the spatial and pitch domains. The time step was varied from $\Delta t=0.2$ (blue curve), to $\Delta t=0.1$ (orange curve), and $\Delta t=0.05$ (green curve). The tokamak was assumed to have a minor radius of $a=0.5\;[\text{m}]$, an inverse aspect ratio of $a/R_{0}=1/3$, a magnetic field of $B_{0}=2\;[\text{T}]$, and a constant safety factor $q=2$.
Aside from the speed $v$, this reduced set of guiding center equations conserves toroidal canonical momentum in the absence of collisions which can be written as:
$p_{\varphi}=m_{i}R\frac{B_{\varphi}}{B}v_{\|}-Ze\psi,$ (21)
For circular flux surfaces, and making analogous approximations as the guiding center equations [Eq. (15)], allows the normalized toroidal canonical momentum to be expressed as:
$\frac{p_{\varphi}}{R_{0}m_{i}v_{Ti}}=\frac{R}{R_{0}}\xi v-\frac{1}{2}\frac{1}{\rho^{*}_{i}}\frac{a}{R_{0}}\frac{r^{2}}{q},$ (22)
where we have taken the safety factor to be a constant for simplicity. Figure 1 provides a numerical demonstration of the conservation of toroidal canonical momentum conservation for ions with 20 and 50 keV. Here, while the error in toroidal canonical momentum conservation grows with time, after a time $2\times 10^{6}a/v_{Ti}$, the error remains below $10^{-2}$ for a time step of $\Delta t=0.1a/v_{Ti}$, for 20 keV ions, and $\Delta t=0.05a/v_{Ti}$, for 50 keV ions. An error of less than one percent will be sufficiently accurate for the present study. JONTA simulations described in the remainder of this work will use a time step of $\Delta t=0.1a/v_{Ti}$ when studying 20 keV ions, and $\Delta t=0.05a/v_{Ti}$ for 50 keV ions.
The collision operator is taken to be Lorentz operator [Eq. (2)], where the deflection frequency in the limit of $v\gg v_{Ti}$ can be expressed as:
$\frac{a\nu_{D}}{v_{Ti}}=\left(a\frac{\hat{\nu}_{si}}{v_{Ti}}\right)\frac{v^{3}_{Ti}}{v^{3}}=\pi\left(n_{i}r^{2}_{e}a\right)Z^{2}_{s}Z^{2}_{i}\left(\frac{m_{e}c^{2}}{T_{i}}\right)^{2}\ln\Lambda\frac{v^{3}_{Ti}}{v^{3}},$
where $r_{e}=2.8179\times 10^{-15}\;[\text{m}]$ is the classical electron radius, $Z_{i}$ is the charge state of the background ions, and $Z_{s}$ is the charge state of the energetic ion. The Lorentz collision operator is implemented by the Monte Carlo equivalent Boozer and Kuo-Petravic (1981)
$\xi_{n+1}=\xi_{n}\left(1-\nu_{D}\Delta t\right)\pm\sqrt{\left(1-\xi^{2}_{n}\right)\nu_{D}\Delta t},$ (23)
where $\Delta t$ is the collisional time step, which is taken to be the same as the collisionless time step. While the Lorentz collision operator does not account for the slowing down of energetic ions, a critical aspect of describing energetic particle evolution, it does drive cross field transport. This will thus allow us to identify energetic ions that are lost rapidly, versus those that will be much better confined. The addition of a more complete set of physical processes including 3D magnetic fields, energetic particle modes, and collisional slowing down will be pursued in future work.
III.3 Example Ion Orbits and Coordinate Conventions
Figure 2: Circular flux surface geometry used in this work and example collisionless ion orbits. Panel (a) shows co-current ions with a passing ion with initial pitch $\xi=0.8$ (red curve), and a trapped ion with initial pitch $\xi=0.3$ (green curve). Panel (b) shows counter-current ions with a passing ion with initial pitch $\xi=-0.8$ (red curve), and a trapped ion with initial pitch $\xi=-0.3$ (green curve). Flux surface contours are shown in blue, with the example orbits shown in red. The black marker indicates the initial position of the ions, and arrows indicate their direction. The grad-B drift points downward in our example geometry. A deuterium ion with $20\;\text{keV}$, a tokamak with a minor radius of $0.5\;\text{m}$, inverse aspect ratio $a/R_{0}=1/3$, constant $q$-profile of $q=2$, and an on-axis magnetic field strength of $B_{0}=2\;\text{T}$ was assumed.
Before discussing solutions to the adjoint drift kinetic equation, it will be useful to describe the magnetic geometry employed. A plot of the circular flux surfaces assumed together with four example ion orbits are shown in Fig. 2. Here, the blue contours indicate magnetic flux surfaces, and the red and green curves indicate example passing and trapped ion orbits, respectively. The toroidal magnetic field and current are assumed to come out of the page, with the poloidal magnetic field in the counter-clockwise direction. Passing ions with a positive pitch (co-current), will thus rotate in the counter-clockwise direction when their motion is projected onto the poloidal plane as they propagate along a magnetic field line, with a grad-B drift vertically downward. Ions with a negative pitch (counter-current) will move in a clockwise sense when moving along field lines, but will still have a vertically downward grad-B drift. From Fig. 2, it is apparent that co-current ions [Fig. 2(a)] born on the weak field side will tend to be better confined compared to counter-current ions [Fig. 2(b)] born at the same location, due to the magnetic drifts taking the counter-current ions closer to the spatial boundary. Furthermore, a counter-current propagating passing ion when scattered into the trapped region will move onto an orbit that takes it closer to the plasma edge, as evident from Fig. 2(b), such that these ions will be more susceptible to being lost.
IV Escape Time
Figure 3: Loss histories for 20 keV (panel a.) and 50 keV (panel b.) ions. Solid lines indicate the training loss whereas ‘x’ markers indicate the test loss. Six million training points were used and two million test points.
Figure 4: (a) Escape time for an ion initialized in the counter-current direction with a pitch of $\xi=-0.8$, after different stages of training. The profile of the ion’s escape time is shown in the panels (a), (b) and (c), with the residual indicated in panels (d), (e) and (f). The first column is after the SOAP phase of training (50,000 epochs), the second column is 10,000 into the SSBroyden phase (60,000 net epochs), and the final column is after 50,000 epochs of SSBroyden (100,000 epochs total). The quantity plotted is $\text{log}_{10}\left(1+T\right)$, yielding a $\text{log}_{10}$ scale for large values of $T$, but vanishing when $T=0$. A deuterium ion with $10\;\text{keV}$ was assumed.
This section will describe solutions for the escape time for a fast ion population at 20 and 50 keV utilizing the PINN implementation described in Sec. III.1, along with a comparison to solutions evaluated with JONTA. Loss histories for ions with 20 and 50 keV are shown in Fig. 3. For these cases, we assumed a midsize tokamak with a minor radius of $a=0.5\;\text{m}$, aspect ratio $R_{0}/a=3$, and deuterium fast ions. The background deuterium plasma is assumed to have a temperature of $1\;\text{keV}$, a density of $n_{i}=10^{20}\;\text{m}^{-3}$, and the safety factor is taken to be $q=2$. A representative Coulomb logarithm of $\ln\Lambda=15$ was used in the collision operator. Each of these quantities were taken to be constant across the plasma radius, though spatial profiles could easily be introduced. A fully connected feedforward network with seven layers each with a width of fifty neurons is used. The SOAP optimizer is used for the first 50,000 epochs, with SSBroyden used for the remaining 25,000 epochs. From Fig. 3, it is evident that SOAP drives a slow, but steady decay of the loss, with SSBroyden driving a sharper decline. We note, however, that the SOAP optimizer being employed in this paper is able to effectively use the Blackwell 200 GPU employed in this work, whereas the SciPy implementation of SSBroyden primarily uses a CPU. As a result, the vast majority of the training time is spent during the SSBroyden phase of optimization.
Considering the evolution of the predicted solution during training, Fig. 4(a) shows the predicted escape time for a 20 keV ion after the SOAP optimization phase, 10,000 epochs into the SSBroyden phase [Fig 4(b)], and the final predicted escape time [Fig 4(c)]. A reference solution from JONTA is shown in Fig. 9(d) below for comparison. It is evident that SOAP is able to capture the structure of the solution near the tokamak edge, but fails to recover the long confinement time of ions initially located at small minor radii. This is due to the large discrepancy in time scales between ions that begin on loss orbits, and are thus lost on a transit time scale, and those initially located deep in the plasma interior, which must collisionally diffuse out of the plasma. This latter process takes place on a far longer time scale, roughly $10^{5}$ slower for the present example, which SOAP struggles to recover. After 10,000 epochs with SSBroyden, the training and test loss of the residual and boundary terms drop further, and it is evident from Fig. 4 that the relatively long confinement time of ions initially located in the interior is better captured, though its value is still substantially under predicted. Finally, after 25,000 epochs of SSBroyden, the predicted escape time of ions initially located in the interior slowly increases, though the final value is still below that predicted by the particle-based code JONTA.
Figure 5: Test loss distribution for 20 keV ions. A uniform random distribution was used, where the size of the markers is proportional to the magnitude of the residual.
Turning to the distribution of training losses, six million training points were used with two million uniformly distributed test points applied. The training points were taken to be distributed by a Hammersley sequence. Two dimensional projections of the resulting test error are shown in Fig. 5, with the magnitude of the markers set to be proportional to the magnitude of the residual. Hence, darker regions represent regions with the largest residuals. From Fig. 5, the outer thirty percent of the tokamak contained the largest losses. This follows due to this region containing direct loss orbits, whose detailed form can be difficult to resolve. It is also evident that the outer five percent of the plasma radius has somewhat smaller residuals due to the concentration of training points in this region. Finally, regions with $r>0.99$ and poloidal angles on either the outboard ($\Delta\theta<\theta<\Delta\theta$) or inboard ($\theta>-\pi+\Delta\theta$ and $\theta<\pi-\Delta\theta$) midplanes where set to zero, leading to empty regions in Fig. 5(a). For this example we set $\Delta\theta=0.1$. As described further below, these regions contain discontinuities in the solution arising from changes in boundary conditions between the lower and upper half of the tokamak, which cannot be precisely resolved. By removing training points from this region, this allows the PINN to find a solution that smoothly transitions across these narrow regions.
The phase space distribution of fast ion escape times can be conveniently illustrated by taking slices of the $\left(R,Z,\xi\right)$ of the PINN’s solution. The escape time of ions born with $\theta=0$ is shown in Fig. 6. Here, ions in the edge of the plasma can be lost quickly due to direct orbit loss, with ions initially located deeper in the plasma having much longer confinement times. Further noting that the white contour indicates the trapped-passing boundary, it is evident that co-current passing ions (positive pitches above the white contour) have far better confinement in comparison to counter-current ions. Similarly, in the trapped region, ions that begin with $\xi>0$ are much better confined compared to ions with $\xi<0$. Comparing the left and right panels of Fig. 2, this difference in confinement results from trapped ions born with $\xi<0$ at $\theta=0$ having orbits that extend to larger radii [Fig. 2(b)], in contrast to trapped ions born with $\xi>0$ [Fig. 2(a)]. The ions with the worst confinement are those just inside the trapped-passing boundary and $\xi<0$, which are expected to have the largest banana width, and thus be most susceptible to orbit loss.
Figure 6: (a) Escape time for an ion initially located at $\theta=0$. The quantity plotted is $\text{log}_{10}\left(1+T\right)$, yielding a $\text{log}_{10}$ scale for large values of $T$, but vanishing when $T=0$. The white curve indicates the trapped-passing boundary. (b) Residual of Eq. (6). A deuterium ion with $20\;\text{keV}$ was assumed.
Figure 7: Slices of escape time in units of $\log_{10}(1+T)$ for different initial pitch values. Panel (a) is for $\xi=0.8$, (b) is for $\xi=0.3$, (c) is for $\xi=-0.3$, and (d) is for $\xi=-0.8$. The solid white curve indicates the location of the trapped-passing boundary, where ions born with major radii $R$ to the right of the white curve are trapped.
Figure 7 shows time slices of major radius $R$ and vertical height $Z$ for different pitches. These slices allow for the difference in confinement of co and counter current ions to be investigated more clearly. Considering passing co-current ions ($\xi=0.8$), these ions will propagate in the counter-clockwise direction in Fig. 7(a). Noting that ion orbits drift downward, this will result in ions that are initially on the outboard side not being on direct loss orbits, in contrast to those born on the inboard side. This leads to a notable in-out asymmetry of the fast ion population, as evident in Fig. 7(a). This trend is reversed for passing ions that begin in the counter-current direction [Fig. 7(d)]. A more subtle feature is that counter-current ions that are not on direct loss orbits are more likely to be lost in comparison to co-current ions. This is evident by the orange region in Fig. 7(d). The relatively poor confinement of these counter-current ions is due to the collisional trapping of these ions results in the orbit exploring larger minor radii [compare the red and green curves in Fig. 2(b)]. Hence, passing counter-current ions that are within a banana width of plasma edge are far more susceptible to being lost to the plasma compared to passing co-current ions.
Figure 8: Slices of escape time in units of $\log_{10}(1+T)$ for different initial radii. Panel (a) is for $r/a=0.999$, (b) is for $r/a=0.95$, (c) is for $r/a=0.9$, and (d) is for $r/a=0.8$. The while curves are the trapped-passing boundary.
Cross-cuts of the pitch and poloidal angle distribution of loss times at different radii are shown in Fig. 8. Here, when approaching $r/a=1$ a discontinuity appears in the solution. This is evident in Fig. 8(a), where for $\theta\approx 0$ and $\theta\approx\pi$, the solution changes rapidly when traversing these discontinuous regions. The origin of this discontinuity is that for ions in the lower half of the tokamak (i.e. $\theta<0$) and $r/a=1$, they are lost immediately. In contrast, ions located just above $\theta=0$ or $\theta=\pi$ are often not on direct loss orbits, and thus can be confined for a long time. For example, co-current ions just above $\theta=0$, will drift inward, and thus have a finite escape time, whereas those just below $\theta=0$ are lost immediately to the boundary. An analogous discontinuity in the solution is present for $\theta=\pi$. To relax these discontinuities, we have removed all training and test points from these regions, such that the PINN is able approximate these discontinuities with a sharply varying solution, as evident in Fig. 8(a). Moving radially away from the $r/a=1$ boundary, panels (b)-(d) of Fig. 8 show the solution at radial locations between $r/a=0.95$ and $r/a=0.8$, indicating vastly different escape times of ions initially located at different pitch and poloidal angles. The trapped-passing boundary is also indicated in this figures, with the escape time of the solution often varying sharply across the trapped-passing region.
Figure 9: Slices of escape time in units of $\log_{10}(1+T)$ for different initial pitch values and an energy of $20\;\text{keV}$ computed from the JONTA code. Panel (a) is for $\xi=0.8$, (b) is for $\xi=0.3$, (c) is for $\xi=-0.3$, and (d) is for $\xi=-0.8$. Ten million markers were used, where the particles were integrated for $t_{final}=2\times 10^{6}$.
The escape time can also be computed by the particle-based solve JONTA described in Sec. III.2. Here, marker particles are initialized randomly across the spatial domain at a given pitch. Once a particle’s minor radius exceeds $r_{wall}$, its escape time and location is saved. By pushing a large number of marker particles, the average escape time of the marker particle can be computed. Noting the conceptual simplicity of this approach, we will use it to validate the predictions of the PINN. Figure 9 shows four slices in the $\left(R,Z\right)$ plane at constant pitch computed from JONTA. Here, ten million marker particles are randomly distributed at a given pitch, with their loss times recorded after they escape from the domain. These escape times are then binned, allowing for the average escape time to be computed over the $\left(R,Z\right)$ plane for a given pitch. The particles were evolved until $t_{final}=2\times 10^{6}$, a time at which less than 0.001 percent of the original particles remained in the domain. Four cross cuts are shown in Fig. 9. These four cross cuts are for the same physics parameters and pitches as Fig. 7 above. Good agreement between the average escape time predicted by JONTA and the PINN are evident. In particular, both approaches are able to recover the rapid escape of ions that are initiated on direct loss orbits (blue regions), ions lost due collision trapping/detrapping directly onto a loss orbit (yellow and green regions), and ions initially located near the magnetic origin that must slowly diffuse out (red regions). The JONTA simulations also reveal a limitation of our implementation of a PINN. In particular, while the PINN is able to distinguish the complex edge structure of confined versus unconfined orbits, with excellent agreement with predictions of JONTA, it struggles to accurately capture the very long escape time for ions initially located deep inside the plasma. Specifically, the dark red region evident in Fig. 9 of the JONTA data, is not present in the PINN’s solution (see Fig. 7), indicating roughly a factor of two discrepancy in predicted confinement for these 20 keV well confined ions. The reason for this quantitative inaccuracy is that the PINN struggles to simultaneously resolve the fast direct orbit dynamics together with the slow collisional transport of the core ions, though as indicated in Fig. 4, this discrepancy slowly shrinks as the model is trained. The resolution of this disparate physics will require either the inclusion of data in training the PINN, using data taken directly from JONTA for example, or modifications to the training of the PINN. These extensions will be discussed further below.
Figure 10: Comparison of escape time in units of $\log_{10}(1+T)$ predictions from the PINN (top row) and JONTA (bottom row). Ions were assumed to have an initial energy of $50\;\text{keV}$. Panels (a,e) are for $\xi=0.8$, (b,f) are for $\xi=0.3$, (c,g) are for $\xi=-0.3$, and (d,h) are for $\xi=-0.8$. Ten million markers were used in the JONTA simulations, where the particles were integrated for $t_{final}=10^{7}$.
Now considering an energetic ion population with $50\;\text{keV}$, similar physical trends are evident, but the larger ion orbits, lead to a larger number of ions beginning on direct loss orbits. In addition, the higher energy of the particles leads to a lower collision time, and thus energetic ions initially located in the core have a longer escape time. This latter property results in the problem exhibiting greater stiffness, with a faster bounce time but slower collision time, resulting in the PINN struggling to accurately resolve the escape time of the best confined orbits. This is evident in Fig. 10, where cross sections in the $\left(R,Z\right)$ plane computed both with JONTA and the PINN are compared. While the PINN fails to capture the precise escape time of the best confined particles, it does succeed in accurately capturing the phase space structure of the escape time. It thus acts as an effective tool for delineating well confined regions of phase space, from those where ions are directly lost, or lost after a small number of collisional scattering events. We note that while for the case of $20\;\text{keV}$ ions roughly a factor of two difference in the escape time was present for the best confined ions, for $50\;\text{keV}$ nearly an order of magnitude difference in the prediction of the best confined orbits is evident. This larger deviation in the prediction of the best confined ions is due to the increased time scale separation between the collisionless ion orbits and the slow collision time.
V Conclusions
The inhomogeneous adjoint of the drift kinetic equation was used to evaluate the escape time of energetic ion populations. This quantity, computed in the limit of an axisymmetric tokamak plasma, captures fast ion losses due to direct orbit loss and collisional transport, thus providing a metric for energetic particle confinement. A PINN was used to evaluate the phase space dependence of the escape time, a challenging task due to the strong time scale separation between the fast bounce time of energetic ions, and their slow collision rate. While the offline training time of the PINN is substantial, the online inference time is rapid, typically microseconds per prediction. Although the PINN failed to quantitatively capture the escape time of the best confined ion orbits, it is able to robustly recover the phase space structure of the escape time, and thus acts as a means of distinguishing well confined orbits versus those on direct loss orbits, or those that are directly lost after being scattered into or out of the trapped or passing regions. We also note that the present analysis did not include fast ion slowing down, where the slowing down time for the 20 and 50 keV ions considered in this work would be far shorter than the escape times of the best confined orbits.
For this proof-of-principle demonstration of the proposed framework we considered several simplifications to the complete system. Perhaps the primary simplifications were the assumption of (i) an axisymmetric magnetic geometry, (ii) a simplified collision operator, and (iii) idealized plasma profiles and shaping. Regarding item (i), since PINNs to date have not been used to solve the drift kinetic equation, beginning with an axisymmetry geometry was a logical starting point for assessing this approach’s strengths and weaknesses. We anticipate extending the present approach in future work to 3D magnetic geometry, though such an extension will likely require incorporating particle data into the training of the neural network. The particle-based solver JONTA will provide a convenient tool through which such data can be generated, where by providing an accurate calculation of the escape time at specific phase space locations, these may be included in the first term of the loss function defined by Eq. (10). By providing high fidelity data points at internal locations, for example in regions with large residuals or for the best confined ion orbits where the present implementation of a PINN has struggled, this will likely expedite the training of the PINN while providing a systematic means of improving its accuracy. Our aim will be to leverage data as a means of both treating more challenging scenarios, such as non-axisymmetric geometry, along with facilitating the learning of a parametric solution to the drift kinetic equation, enabling the escape time to be evaluated for a broad range of plasma conditions.
An additional simplification was the use of a Lorentz collision operator, which does not account for the slowing down of energetic particles, a quantity of critical interest when assessing the amount of energy deposited by energetic ions into the bulk plasma. Despite this limitation, including Lorentz collisions allows for the collisional transport of fast ions to be assessed. Thus, while not providing an accurate indication of the escape time for well confined energetic particles, since such particles will likely slow down substantially during their long confinement time, it does provide a means of rapidly assessing particles on direct loss orbits, or those that are directly lost after being collisionally scattered into or out of the trapped particle region. The predictions of the present work will thus be most relevant to the outer region of the tokamak, where energetic particles are lost on relatively short time scales. Finally, homogeneous plasma profiles and a circularly shaped plasma were assumed in the present work. This is done for convenience, where we don’t anticipate including more realistic plasma profiles to pose a substantial challenge. Furthermore, since PINNs are mesh free, and thus well suited to treating shaped magnetic geometry van Milligen et al. (1995); Kaltsas and Throumoulopoulos (2022); Jang et al. (2024), we do not expect the treatment of shaped axisymmetric geometry to pose a major hurdle.
Due to the strong time scale separation, evaluating fast ion transport is computationally intensive, creating a bottleneck in many optimization workflows. For the JONTA simulations presented in this paper, the shortest simulation required one and a half hours on a single NVIDIA Blackwell 200 GPU, with the longest requiring slightly over a day. Noting that each JONTA simulation was for a fixed value of pitch, computing four distinct pitches (Figs. 9 and 10) thus required several hours to days of computing time. A PINN, which encompasses all values of pitch, and thus only a single PINN must be trained, required approximately two days to train, where approximate results obtained after the SOAP phase of training typically take a few hours. Both approaches are thus computationally intensive, where while the PINN is able to recover smooth predictions of the escape time, it is unable to accurately compute the best confined ions. JONTA, in contrast, requires roughly ten million markers to obtain a relatively smooth, well resolved solution. We note that while the PINN also used a single NVIDIA Blackwell 200 GPU, the implementation of the SSBroyden optimizer does not make efficient use of the GPU, and thus was mostly trained using a CPU, which was the primary bottleneck in the training of the PINN.
We also note that while the present paper has focused on the escape time of energetic ions several straightforward extensions are possible. One such extension would be to treat a broader range of fast ion metrics. Metrics of interest include the probability that an ion impacts a specific region of the wall from a given initial phase space location Särkimäki (2020), or the slowing down probability, which corresponds to the probability that an ion slows down to the thermal energy. Both of these fast ion metrics can be straightforwardly constructed from the steady state homogeneous adjoint equation by applying appropriate boundary conditions. The evaluation of this broader class of fast ion metrics will be left for future work. Finally, while we have considered energetic ions in the present formulation, extensions to the relativistic drift kinetic equation describing runaway electron (RE) evolution and transport can be straightforwardly developed. Such extensions will allow for ‘neoclassical’ corrections to RE generation rates to be inferred Rosenbluth and Putvinski (1997); McDevitt and Tang (2019); Arnaud and McDevitt (2024), or account for spatial transport due to processes such as the Ware pinch Nilsson et al. (2015); McDevitt et al. (2019). While we anticipate that the extremely short bounce time of REs compared to their long collision times will pose a substantial challenge, we note that the impact of ‘neoclassical’ effects on RE generation rates are most important near the critical energy for electrons to run away McDevitt and Tang (2019); Arnaud and McDevitt (2024), an energy that often ranges between several keV to tens of keV during a tokamak disruption. At these modest energies, the time scale separation between the transit or bounce time of a RE and the collision time will be drastically reduced. The extension of the present framework to REs will be pursued in future work.
Acknowledgements.
This work was supported by the Department of Energy, Office of Fusion Energy Sciences under awards DE-SC0024634 and DE-SC0024649, and the University of Florida Division of Sponsored Research Projects. The authors acknowledge the University of Florida Research Computing for providing computational resources that have contributed to the research results reported in this publication.